OpenStudy (anonymous):

sin(4x)=8 sin(x)cos^3(x)-4 sin(x)cos(x) Proofs are beyond confusing...

OpenStudy (freckles):

sounds like you are trying to show this is an identity

OpenStudy (anonymous):

Which I am trying to figure out now lol

OpenStudy (anonymous):

it looks like i should be using 2 sin theta and cos theta

OpenStudy (freckles):

is the above written correctly?

OpenStudy (freckles):

seems like there could be a cube or something missing

OpenStudy (anonymous):

yeah it suppose to be cosine cubed which i have there

OpenStudy (freckles):

well see if you can continue from here \[8 \sin(x) \cos^3(x)-4 \sin(x)\cos(x) \\ \text{ I would start by factoring out } 4 \sin(x)\cos(x) \\ 4 \sin(x) \cos(x)[\cos^2(x)-1] \\ 2 \cdot 2 \sin(x)\cos(x)[-1(1-\cos^2(x))] \]

OpenStudy (freckles):

like right here I see a couple of identities you can use

OpenStudy (freckles):

oops missed a 2

OpenStudy (freckles):

\[8 \sin(x) \cos^3(x)-4 \sin(x)\cos(x) \\ \text{ I would start by factoring out } 4 \sin(x)\cos(x) \\ 4 \sin(x) \cos(x)[2\cos^2(x)-1] \\ 2 \cdot 2 \sin(x)\cos(x)[-1(1-2 \cos^2(x))] \]

OpenStudy (freckles):

beautiful :)

OpenStudy (freckles):

for example guess what 2sin(x)cos(x) or guess what 2cos^2(x)-1 or if you like -1(1-2cos^2(x)) equals

OpenStudy (freckles):

http://www.eeweb.com/tools/math-sheets/images/trigonometry-laws-and-identities-small.png look under double angle identities if you do not recall

OpenStudy (anonymous):

This is what I did

OpenStudy (anonymous):

so is this right?

OpenStudy (freckles):

I like the last 5 lines and the first 2 lines looks like you factored as I did above just didn't end the ( parenthesis thing on the second line

OpenStudy (freckles):

or on the first line

OpenStudy (freckles):

\[8 \sin(x) \cos^3(x)-4 \sin(x)\cos(x) \\ \text{ I would start by factoring out } 4 \sin(x)\cos(x) \\ 4 \sin(x) \cos(x)[2\cos^2(x)-1] \\ \] but yeah that is essentially what I have above

OpenStudy (freckles):

you just need some ( ) in there

OpenStudy (freckles):

this is exactly what I would do to fix your ( ) issue: \[4 \sin(x)\cos(x)(2\cos^2(x)-1) \text{ for first line; just drop that one \between the } \\ 4 \text{ and } \sin(x) \\ \\\]

OpenStudy (freckles):

the second line ... I would change that first ( to a * and then also put ( ) around the 2cos^2(x)-1 like so: \[2 \cdot 2 \sin(x)\cos(x)(2 \cos^2(x)-1)\]

OpenStudy (anonymous):

Thanks soo much again!! I swear so much work under one problem lol @freckles